a projectile is fired with an initial velocity of 60 m/s upward at an angle of 30 degree to the horizontal from a point 80m above a level plain. what horizontal distance will it cover before it strikes the plain?


1
Expert's answer
2022-01-31T07:35:14-0500

Answer


This is vertical speed component


u=60sin 30°=30m/s


Height h=80m


Using second equation of motion


S=ut+0.5at2

Putting all values


-80=30t-4.9t2

So this quadratic equation in t


So it's solutions

t2 -6.12t-16.32=0

This is quadratic equation in t it's solutions are

"t=\frac{6.12±\sqrt{37.45+65.28}}{2}\\=\frac{6.12±\sqrt{102.73}}{2}\\=\frac{6.12±10.14}{2}"

By using positive sign then time

t=10. 26sec

So horizontal distance

D=60 cos 30° *10.26=533.12m